# The Galactic population of canonical pulsars II

### v2

@@ 1 Introduction
Pulsars, discovered by Jocelyn Bell Burnell in 1967 (Hewish et al., 1968) are rotating neutron stars born in a core-collapse supernova. They are highly magnetized and surrounded by a plasma-filled magnetosphere emitting regular pulses of radiation at their spin frequency. Due to the magneto-dipole losses, they lose rotational kinetic energy and their spin period increases. Extensive radio surveys, performed for instance by the Parkes and Arecibo radio-telescopes, so far 3700 pulsars have been discovered (Manchester et al. 2005, see Table 1 listing the canonical pulsars). Although pulsars were first observed in radio, they were later found to be also bright in X-rays, optical and $\gamma$-rays. The Large Area Telescope (LAT) on board of the Fermi satellite, has discovered dozens of radio-quiet $\gamma$-ray pulsars as well as millisecond pulsars (MSPs). Since its launch, the LAT has detected about 300 pulsars (Abdo et al., 2013; Smith et al., 2019, 2023) and unveiled a new perspective, expanding our understanding about pulsar emission mechanisms by increasing the sample of neutron stars detectable through their high energy emission. Moreover, the distinct radiation process involved in high-energy photon production and its unique beaming properties offer an alternative viewpoint on the global pulsar population.
Table 1: Number of known canonical pulsars, with spindown luminosity $\dot{E}$ above and below $10^{28}$ W, and above $10^{31}$ W. The quantities $N_{\rm r}$, $N_{\rm g}$, and $N_{\rm rg}$ are the number of radio-only, gamma only, and radio-loud $\gamma$-ray pulsars, respectively. It should be noted that we excluded the binary pulsars, the pulsars in the Magellanic clouds and in globular clusters. The data have been taken from the ATNF catalogue from https://www.atnf.csiro.au/research/pulsar/psrcat/
@@
Most PPS studies assume that neutron stars are rotating in vacuum (Faucher-Giguère & Kaspi, 2006; Popov et al., 2010; Johnston et al., 2020) or take only radio or $\gamma$-ray emission into account separately (Watters & Romani, 2011; Gullon et al., 2014), or assume a constant magnetic field during the evolution process (Gonthier et al., 2002; Faucher-Giguère & Kaspi, 2006; Johnston & Karastergiou, 2017). Dirson et al. (2022) have modelled for the first time the population of both the $\gamma$-ray and radio pulsars, taking into account the state-of-the-art force-free magnetosphere model in conjunction with a prescription for the magnetic field decay. The neutron star radiation is produced by relativistic charged particles flowing within their magnetosphere. Therefore the plasma back reaction must be taken into account in the evolution of the pulsar period and magnetic inclination angle. Whereas the location of the $\gamma$-ray emission site is still under debate, different regions have been suggested. These are for instance the polar cap region (Sturrock, 1971; Ruderman & Sutherland, 1975; Daugherty & Harding, 1982, 1996) at low altitudes, the slot gap along the last open magnetic field lines (Arons, 1983; Muslimov & Harding, 2004; Harding et al., 2008; Harding & Muslimov, 2011), or the outer gaps at high altitudes within the light-cylinder, in the outer magnetosphere (Cheng et al., 1986; Hirotani, 2008; Takata et al., 2011). Emission is also possible outside the light cylinder, in the so-called striped wind (Pétri, 2009, 2011) and was used for the first time by Dirson et al. (2022) in a PPS study. The radio emission is described as usual by the polar cap model. In this work as in Dirson et al. (2022) only the canonical pulsar population in our Galaxy is reproduced. The canonical pulsar population is defined by isolated pulsars which are neither magnetars nor millisecond pulsars.
@@ 2.1 Generation of pulsars
$p(\log(P_{0}))=\frac{1}{\sigma_{p}\sqrt{2\pi}}e^{-(\log P_{0}-\log\bar{P})^{2}/(2\sigma_{p}^{2})}\ .$ (4)
@@ 2.2 Pulsar evolution
where $\alpha_{d}$ is a constant parameter controlling the rate of the magnetic field decay and $\tau_{d}$ the typical decay timescale, which depends on the initial magnetic field as demonstrated by magneto-thermal evolution models (Viganò et al., 2013). We took these results into account by defining the decay timescale $\tau_{1}$ for a magnetic field strength $B_{1}$ such that $\tau_{1}\,B_{1}^{\alpha_{d}}=\tau_{d}\,B_{0}^{\alpha_{d}}$. Thus, $\tau_{1}$ is another magnetic field decay time scale, corresponding to a certain magnetic field value $B_{1}$. Three different $\tau_{1}$ are considered in this work and are randomly chosen for each pulsar generated : $1.5\times 10^{5}$ yr or $3.5\times 10^{5}$ yr or $2.5\times 10^{6}$ yr with each of these timescales corresponds to a magnetic field value $B_{1}$ of $1\times 10^{8}$ T, $3\times 10^{8}$ T and $2\times 10^{9}$ T (these values of $\tau_{1}$ and $B_{1}$ are taken from Fig. 10 of the work of Viganò et al. (2013)). Then the probabilities to get these values of $\tau_{1}$ and $B_{1}$ to be adopted for one pulsar are 0.23, 0.46 and 0.31 respectively (these probabilities were chosen as best educated guesses after several run with different values). The value of $\alpha_{d}$ can be found in Table 5 of Sect. 3.
@@ 2.4 The death line
Several equations for the death line have been proposed in the literature (Chen & Ruderman, 1993; Zhang et al., 2000; Gil & Mitra, 2001; Mitra et al., 2019). The most suitable one in our study is that of Mitra et al. (2019) where they found an expression relating $P$ and $\dot{P}$ by
@@ 2.5.1 Radio detection model
where $h_{\rm em}$ is the emission height, $P$ is the spin period of the pulsar and $c$ is the speed of light. The emission height is taken constant with an average value of $h_{\rm em}=3.10^{5}$ m, estimated from observations of a sample of pulsars by Weltevrede & Johnston (2008); Mitra (2017); Johnston & Karastergiou (2019); Johnston et al. (2023). The cone half-opening angle $\rho$ estimated in eq. (25) holds only for the last open field lines of a magnetic dipole and can only be applied for slow pulsars where the radio emission altitude is high enough for the multipolar components to decrease significantly and become negligible.
@@
$\tau_{DM}=\frac{e^{2}\Delta f_{ch}\,DM}{\pi\,m_{e}\,c\,f^{3}}$ (30)
$e$ is the electron charge, $m_{e}$ its mass, $\Delta f_{ch}$ the width of frequency of the instrument channel, $f$ the observing frequency and $DM$ the dispersion measure. Furthermore, in order to determine the influence of scattering by an inhomogeneous and turbulent ISM through $\tau_{scat}$, the empirical fit relationship from Krishnakumar et al. (2015) is used
@@
Both $\tau_{DM}$ and $\tau_{scat}$ depend on the dispersion measure, which is found for each pulsar by running the code of Yao et al. (2017) converting the distance of a pulsar into the dispersion measure thanks to their state-of-the art model of the Galactic electron density distribution. The radio survey parameters used are the ones from the Parkes Multibeam Pulsar Survey (PMPS) taken from Manchester et al. (2001), see Table 4.
@@ 3.1.1 Without a death line
In order to verify whether there is a correlation between old pulsars and those with a high spin period or not, Fig. 2 shows in green pulsars with a characteristic age greater than $10^{8}$ yr, corresponding to the oldest pulsars of the simulation. It coincides with the excess found in Fig. 2. We conclude that canonical pulsars older than $10^{8}$ yr measured as their real age in our simulations should be discarded at the end of the simulation, since the characteristic age is overestimating the real age of pulsars. Indeed canonical pulsars older than this do not exist. The Galactic potential forbids many pulsars to escape from the Milky Way, they remain bound to the Galaxy, in contrast to the older pulsars found in Dirson et al. (2022), that were escaping the Galaxy easily because of the absence of this potential. Fig. 4 shows that many pulsars are in what is commonly called the graveyard of pulsars at the bottom right of the $P-\dot{P}$ diagram, the death valley being shown in Fig. 4 but not implemented.
@@
Given this excess of long period pulsars when neglecting the death line, we decided to implement it, in order to stick closer to reality.
@@ 3.5 $\gamma$-ray detection
Let us now focus on the $\gamma$-ray pulsar population. First, we note that the total number of pulsars detected is significantly larger in our sample (307 detections) compared to the observations (173 detections), when using the Fermi/LAT instrument sensitivity. However, the simulated $\gamma$-ray only and radio loud $\gamma$-ray pulsars are in the expected area of Fig. 20, furthermore, the KS test gives a p-value of 0.52 and $\approx 0$ for $\dot{P}$ and $P$ respectively. The null hypothesis can not be rejected for $\dot{P}$ but it is rejected for $P$. However, when the KS test is conducted on the $\gamma$-ray population only we get p-values of 0.03 and 0.14 for $\dot{P}$ and $P$ respectively, therefore here the null hypothesis is rejected for $\dot{P}$ but not for $P$. While for the radio loud $\gamma$-ray population p-values of 0.11 and $\approx 0$ are obtained for $\dot{P}$ and $P$ respectively. Thus, we thought the low p-value obtained for $P$ for both population was probably because of the 7 high spin period radio loud $\gamma$-ray pulsars seen on Fig. 20, but even when conducting the KS test without these pulsars gave a p-value $\geq 0.05$ for $\dot{P}$ and well below $0.05$ for $P$. In addition, we also compare the $\gamma$-ray flux distribution of the simulated and observed populations. In Fig.21 we notice that both distributions peak at the same flux level around $10^{-13.7}$ W.m-2. Globally, the number of pulsars in the simulation with a flux greater than $10^{-13.4}$ W.m-2 is very close to the observations. However, between $10^{-14.4}$ W.m-2 and $10^{-14}$ W.m-2, there are much more pulsars in the simulation. This bi-modality in the simulated $\gamma$ray flux distribution is an artefact introduced by our abrupt condition on the threshold $F_{\rm min}$ depending on the radio detection of a pulsar and on a sufficiently low galactic latitude. However if we smoothly change $F_{\rm min}$ from $4\times 10^{-15}$ to $16\times 10^{-15}$ (the two $F_{\rm min}$ values used in this work), the bi-modality would almost disappear. The higher number of detected pulsars in the simulations could be related to the unidentified sources in the Fourth Fermi-LAT catalogue of $\gamma$-ray sources (4FGL). Since many unidentified Galactic sources are listed in this catalogue, many $\gamma$-ray pulsars might have been observed without being identified as such yet. Nonetheless, assuming that every unidentified $\gamma$-ray pulsar in this catalogue will be recognized as such by a future more sensitive instrument and that a significant difference between the number of simulated and observed pulsars subsists, this could hint to some shortcomings faced by the striped wind model. Let us mention some of these possible flaws of geometrical or physical nature. First the $\gamma$-ray beam opening angle could be smaller than the one used due to some plasma recollimation effect within the striped wind. This would reduce the number of detected pulsars. Second the luminosity function could also deviate from the prescription adopted in (32). Another dependence on $B$ and $P$ or equivalently on $P$ and $\dot{P}$ could be introduced, maybe adding a third parameter like the cut-off energy to better describe the $\gamma$-ray luminosity. Furthermore, we compare our results with the PPS work of Pierbattista et al. (2012), where they simulated the young $\gamma$-ray pulsar population with 4 different models : Outer Gap (OG), Slot Gap (SG), Polar Cap (PC) and Outer Polar Cap (OPC) models. To summarize their findings, their best model is the OPC model, showing almost identical results as ours, $P-\dot{P}$ shape and number of detection, except that we detect $\gamma$-ray pulsars a little farther away from Earth.
@@ 5 Summary
Nonetheless, simulating 1 million pulsars allowed to show that with or without a death line the results were alike and therefore, in the case where the oldest simulated pulsar is $4.1\times 10^{7}$ yr (its real age), a death line is not necessary to explain the observations. In this paper we showed two approaches to reproduce the observed canonical pulsar population, however when we consider the approach where the oldest pulsar simulated is only $4.1\times 10^{7}$ yr we state that no older pulsars might exist in the Galaxy, while the Milky Way is very much older than this. There is apparently no reason not to simulate pulsars that are older than this, and if there is one we do not know it, they become harder to detect because of their wide pulse profile with their age, but some are still detectable. If this situation is the reality then it means that the death line is not necessary in pulsar population synthesis to explain observations. When the simulation is ran with 10 millions pulsars, the death line becomes a necessity. However this simulation gives results less similar to observations.

### v3

@@ 1 Introduction
Pulsars, discovered by Jocelyn Bell Burnell in 1967 (Hewish et al., 1968) are rotating neutron stars born in a core-collapse supernova. They are highly magnetized and surrounded by a plasma-filled magnetosphere emitting regular pulses of radiation at their spin frequency. Due to the magneto-dipole losses, they lose rotational kinetic energy and their spin period increases. Extensive radio surveys, performed for instance by the Parkes and Arecibo radio-telescopes, so far 3700 pulsars have been discovered (Manchester et al. 2005, see Table 1 listing the canonical pulsar population defined by isolated pulsars with periods longer than 20 msec and that are not magnetars). Although pulsars were first observed in radio, they were later found to be also bright in X-rays, optical and $\gamma$-rays. The Large Area Telescope (LAT) on board of the Fermi satellite, has discovered dozens of radio-quiet $\gamma$-ray pulsars as well as millisecond pulsars (MSPs). Since its launch, the LAT has detected about 300 pulsars (Abdo et al., 2013; Smith et al., 2019, 2023) and unveiled a new perspective, expanding our understanding about pulsar emission mechanisms by increasing the sample of neutron stars detectable through their high energy emission. Moreover, the distinct radiation process involved in high-energy photon production and its unique beaming properties offer an alternative viewpoint on the global pulsar population.
Table 1: Number of known canonical pulsars, with spindown luminosity $\dot{E}$ above and below $10^{28}$ W, and above $10^{31}$ W. The quantities $N_{\rm r}$, $N_{\rm g}$, and $N_{\rm rg}$ are the number of radio-only, gamma only, and radio-loud $\gamma$-ray pulsars, respectively. It should be noted that we excluded the binary pulsars, the pulsars in the Magellanic clouds and in globular clusters, then we selected the pulsars with a magnetic field between $6\times 10^{6}$ T and $4.4\times 10^{9}$ T. The data have been taken from the ATNF catalogue from https://www.atnf.csiro.au/research/pulsar/psrcat/
@@
Most PPS studies assume that neutron stars are rotating in vacuum (Faucher-Giguère & Kaspi, 2006; Popov et al., 2010; Johnston et al., 2020) or take only radio or $\gamma$-ray emission into account separately (Watters & Romani, 2011; Gullon et al., 2014), or assume a constant magnetic field during the evolution process (Gonthier et al., 2002; Faucher-Giguère & Kaspi, 2006; Johnston & Karastergiou, 2017). Dirson et al. (2022) have modelled for the first time the population of both the $\gamma$-ray and radio pulsars, taking into account the state-of-the-art force-free magnetosphere model in conjunction with a prescription for the magnetic field decay. The neutron star radiation is produced by relativistic charged particles flowing within their magnetosphere. Therefore the plasma back reaction must be taken into account in the evolution of the pulsar period and magnetic inclination angle. Whereas the location of the $\gamma$-ray emission site is still under debate, different regions have been suggested. These are for instance the polar cap region (Sturrock, 1971; Ruderman & Sutherland, 1975; Daugherty & Harding, 1982, 1996) at low altitudes, the slot gap along the last open magnetic field lines (Arons, 1983; Muslimov & Harding, 2004; Harding et al., 2008; Harding & Muslimov, 2011), or the outer gaps at high altitudes within the light-cylinder, in the outer magnetosphere (Cheng et al., 1986; Hirotani, 2008; Takata et al., 2011). Emission is also possible outside the light cylinder, in the so-called striped wind (Pétri, 2009, 2011) and was used for the first time by Dirson et al. (2022) in a PPS study. The radio emission is described as usual by the polar cap model. In this work as in Dirson et al. (2022) only the canonical pulsar population in our Galaxy is reproduced.
@@ 2.1 Generation of pulsars
$p(\log(P_{0}))=\frac{1}{\sigma_{p}\sqrt{2\pi}}e^{-(\log P_{0}-\log\bar{P})^{2}/(2\sigma_{p}^{2})}$ (4)
@@ 2.2 Pulsar evolution
where $\alpha_{d}$ is a constant parameter controlling the rate of the magnetic field decay and $\tau_{d}$ the typical decay timescale, which depends on the initial magnetic field as demonstrated by magneto-thermal evolution models (Viganò et al., 2013). We took these results into account by defining the decay timescale $\tau_{1}$ for a magnetic field strength $B_{1}$ such that $\tau_{1}\,B_{1}^{\alpha_{d}}=\tau_{d}\,B_{0}^{\alpha_{d}}$. Thus, $\tau_{1}$ is another magnetic field decay time scale, corresponding to a certain magnetic field value $B_{1}$. Three different $\tau_{1}$ are considered in this work and are randomly chosen for each pulsar generated : $1.5\times 10^{5}$ yr or $3.5\times 10^{5}$ yr or $2.5\times 10^{6}$ yr with each of these timescales that corresponds to a magnetic field value $B_{1}$ of $1\times 10^{8}$ T, $3\times 10^{8}$ T and $2\times 10^{9}$ T. Then the probabilities to get these values of $\tau_{1}$ and $B_{1}$ to be adopted for one pulsar are 0.23, 0.46 and 0.31 respectively. These three probabilities actually approximate the probability density function for $\tau_{1}$ and $B_{1}$, and were already sufficient to reproduce the canonical population. Finding the complete distribution for $\tau_{1}$ and $B_{1}$ would require further investigations not needed in our study. Moreover, the distributions of $\tau_{1}$ and $B_{1}$ would physically represent the whole ensemble of trajectories possible in the $P-\dot{P}$ diagram for canonical pulsars as in Fig. 10 of Viganò et al. (2013). Therefore, each pulsar has three possible evolutionary paths in the simulation, which is an empirical way to reproduce the $P-\dot{P}$ diagram in a better way than having only one evolutionary path. These particular choices for the discrete values were only guided by trial and errors. The value of $\alpha_{d}$ can be found in Table 5 of Sect. 3.
@@ 2.4 The death line
Several equations for the death line have been proposed in the literature (Chen & Ruderman, 1993; Zhang et al., 2000; Gil & Mitra, 2001; Mitra et al., 2020). The most suitable one in our study is that of Mitra et al. (2020) where they found an expression relating $P$ and $\dot{P}$ by
@@ 2.5.1 Radio detection model
where $h_{\rm em}$ is the emission height, $P$ is the spin period of the pulsar and $c$ is the speed of light. The emission height is taken constant with an average value of $h_{\rm em}=3\times 10^{5}$ m, estimated from observations of a sample of pulsars by Weltevrede & Johnston (2008); Mitra (2017); Johnston & Karastergiou (2019); Johnston et al. (2023). The cone half-opening angle $\rho$ estimated in eq. (25) holds only for the last open field lines of a magnetic dipole and can only be applied for slow pulsars where the radio emission altitude is high enough for the multipolar components to decrease significantly and become negligible.
@@
$\tau_{DM}=\frac{e^{2}\Delta f_{ch}\,DM}{4\,\pi^{2}\,\epsilon_{0}\,m_{e}\,c\,f^{3}}=8.3\times 10^{15}\leavevmode\nobreak\ s\,\frac{\Delta f_{ch}}{f^{3}}\,DM$ (30)
$e$ is the electron charge, $m_{e}$ its mass, $\Delta f_{ch}$ the width of frequency of the instrument channel in Hz, $f$ the observing frequency in Hz and $DM$ the dispersion measure in pc/cm3. Furthermore, in order to determine the influence of scattering by an inhomogeneous and turbulent ISM through $\tau_{scat}$, the empirical fit relationship from Krishnakumar et al. (2015) is used
@@
Both $\tau_{DM}$ and $\tau_{scat}$ are given in units of second (s) and depend on the dispersion measure DM (in units of pc/cm3), which is found for each pulsar by running the code of Yao et al. (2017) converting the distance of a pulsar into the dispersion measure thanks to their state-of-the art model of the Galactic electron density distribution. The radio survey parameters used are the ones from the Parkes Multibeam Pulsar Survey (PMPS) taken from Manchester et al. (2001), see Table 4.
@@ 3.1.1 Without a death line
In order to verify whether there is a correlation between old pulsars and those with a high spin period or not, Fig. 2 shows in green pulsars with a characteristic age greater than $10^{8}$ yr, corresponding to the oldest pulsars of the simulation. It coincides with the excess found in Fig. 2. We conclude that canonical pulsars older than $10^{8}$ yr measured as their real age in our simulations should be discarded at the end of the simulation, since the characteristic age is overestimating the real age of pulsars. Indeed, it appears that we are unable to detect canonical pulsars older than $10^{8}$ yr. The Galactic potential forbids many pulsars to escape from the Milky Way, they remain bound to the Galaxy, in contrast to the older pulsars found in Dirson et al. (2022), that were escaping the Galaxy easily because of the absence of this potential. Fig. 4 shows that many pulsars are in what is commonly called the graveyard of pulsars at the bottom right of the $P-\dot{P}$ diagram, the death valley being shown in Fig. 4 but not implemented. Given this excess of long period pulsars when neglecting the death line, we decided to implement it, in order to stick closer to reality.
@@ 3.5 $\gamma$-ray detection
Let us now focus on the $\gamma$-ray pulsar population. First, we note that the total number of pulsars detected is significantly larger in our sample (307 detections) compared to the observations (173 detections), when using the Fermi/LAT instrument sensitivity. However, the simulated $\gamma$-ray only and radio loud $\gamma$-ray pulsars are in the expected area of Fig. 20, furthermore, the KS test gives a p-value of 0.52 and $\approx 0$ for $\dot{P}$ and $P$ respectively. The null hypothesis can not be rejected for $\dot{P}$ but it is rejected for $P$. However, when the KS test is conducted on the $\gamma$-ray population only we get p-values of 0.03 and 0.14 for $\dot{P}$ and $P$ respectively, therefore here the null hypothesis is rejected for $\dot{P}$ but not for $P$. While for the radio loud $\gamma$-ray population p-values of 0.11 and $\approx 0$ are obtained for $\dot{P}$ and $P$ respectively. Thus, we thought the low p-value obtained for $P$ for both population was probably because of the 7 high spin period radio loud $\gamma$-ray pulsars seen on Fig. 20, but even when conducting the KS test without these pulsars gave a p-value $\geq 0.05$ for $\dot{P}$ and well below $0.05$ for $P$.
In addition, we also compare the $\gamma$-ray flux distribution of the simulated and observed populations. In Fig.21 we notice that both distributions peak at the same flux level around $10^{-13.7}$ W.m-2. Globally, the number of pulsars in the simulation with a flux greater than $10^{-13.4}$ W.m-2 is very close to the observations. However, between $10^{-14.4}$ W.m-2 and $10^{-14}$ W.m-2, there are much more pulsars in the simulation. This bi-modality in the simulated $\gamma$ray flux distribution is an artefact introduced by our abrupt condition on the threshold $F_{\rm min}$ depending on the radio detection of a pulsar and on a sufficiently low galactic latitude. However if we smoothly change $F_{\rm min}$ from $4\times 10^{-15}$ to $16\times 10^{-15}$ (the two $F_{\rm min}$ values used in this work), the bi-modality would almost disappear.
Finally, we compare the simulated and observed $\gamma$-ray light-curve peak separation $\Delta$ in Fig. 22. The separation $\Delta$ is computed according to Pétri (2011) by
$\cos\left(\pi\,\Delta\right)=|\cot\xi\cot\alpha|$ (37)
where $\xi$ is the angle between the line of sight and the rotation axis and $\alpha$ the inclination angle. The distributions are normalized in order to obtain probability distribution functions (p.d.f) for both observations and simulations. The normalization is performed by dividing each distribution function by its total number of pulsars, observed or simulated. Moreover, $\Delta$ is constrained to the interval between 0 and 0.5, because we consider that peak separations $\Delta$ larger than 0.5 can be scaled to $\Delta=1-\Delta$ by a phase shift, since the signal is periodic and represents the phase separation between two peaks and the definition of the first or main peak is arbitrary. Fig. 22 displays a remarkable similarity between the observed and simulated distributions of the $\gamma$-ray peak separation, meaning the striped wind model captures very well this geometric feature.
@@
The higher number of detected pulsars in the simulations could be related to the unidentified sources in the Fourth Fermi-LAT catalogue of $\gamma$-ray sources (4FGL). Since many unidentified Galactic sources are listed in this catalogue, many $\gamma$-ray pulsars might have been observed without being identified as such yet. Nonetheless, assuming that every unidentified $\gamma$-ray pulsar in this catalogue will be recognized as such by a future more sensitive instrument and that a significant difference between the number of simulated and observed pulsars subsists, this could hint to some shortcomings faced by the striped wind model. Let us mention some of these possible flaws of geometrical or physical nature. First the $\gamma$-ray beam opening angle could be smaller than the one used due to some plasma recollimation effect within the striped wind. This would reduce the number of detected pulsars. Second the luminosity function could also deviate from the prescription adopted in (32). Another dependence on $B$ and $P$ or equivalently on $P$ and $\dot{P}$ could be introduced, maybe adding a third parameter like the cut-off energy to better describe the $\gamma$-ray luminosity. Furthermore, we compare our results with the PPS work of Pierbattista et al. (2012), where they simulated the young $\gamma$-ray pulsar population with 4 different models : Outer Gap (OG), Slot Gap (SG), Polar Cap (PC) and Outer Polar Cap (OPC) models. To summarize their findings, their best model is the OPC model, showing almost identical results as ours, $P-\dot{P}$ shape and number of detection, except that we detect $\gamma$-ray pulsars a little farther away from Earth.
@@
Figure 22: Distribution of the $\gamma$-ray peak separation of the simulated population along with the observations from the 3PC catalogue.
@@ 5 Summary
Nonetheless, simulating 1 million pulsars allowed to show that with or without a death line the results were alike and therefore, in the case where the oldest simulated pulsar is $4.1\times 10^{7}$ yr (its real age), a death line is not necessary to explain the observations. In this paper we showed two approaches to reproduce the observed canonical pulsar population, however when we consider the approach where the oldest pulsar simulated is only $4.1\times 10^{7}$ yr we state that no older pulsars might be detected with our current telescopes in the Galaxy, while the Milky Way is very much older than this and therefore has been forming pulsars since its beginning. There is apparently no reason not to simulate pulsars that are older than $10^{8}$ yr, and if there is one we do not know it, they become harder to detect because of their wide pulse profile with their age, but some could still be detectable. If this situation is the reality then it means that the death line is not necessary in pulsar population synthesis to explain observations. When the simulation is ran with 10 millions pulsars, the death line becomes a necessity. However this simulation gives results less similar to observations.